I am having trouble finding the value of b satisfying the description
The line through the points (2,4) and (5,b) with the slope 5/7

Found 2 solutions by richwmiller, rfer:
Answer by richwmiller(17219)

(4-b)/(2-5)=5/7
b = 43/7

Answer by rfer(16322)

(5,6)
graph it using the slope the you can see the (5,6) point